OpenStudy (anonymous):

any help? -6(t-8)-(t+9)=7 t=

OpenStudy (anonymous):

just distribute -6t + 48 - t -9 = 7 combine like terms: -7t + 39 = 7 -39 -7t = -32 t = 32/7